# STable AutoCorrelation Integral Estimator (STACIE): Robust and accurate transport properties from molecular dynamics simulations

### v1

@@ 1 Introduction
[…]
@@
where $F$ is a property-dependent factor. For example, for diffusivity, $F=1$, and for viscosity, $F=V/k_{\mathrm{B}}T$, where $V$ is the volume of the simulation cell, $T$ is the temperature, and $k_{\mathrm{B}}$ is the Boltzmann constant. 18, 3 Note that while it is more common to integrate from 00 to $\infty$, the integral is symmetric, allowing the limits to be extended to $\infty$- ∞ to $\infty$ by including the factor $\frac{1}{2}$. This notation simplifies the discussion of STACIE in later sections. The integration of the ACF also yields the integrated correlation time:
@@
There are several sources of error when numerically integrating the ACF. In the context of MD simulations these include approximations of the function $\hat{x}(t)$, due to force fields, finite-size effects, Ewald convergence, etc. Even if these are well controlled, $\hat{x}(t)$ is a stochastic quantity (due to thermal fluctuations) and can only be sampled at a finite number of time steps (due to limited computational resources). As a result, a direct computation of $\mathcal{I}$ via numerical quadrature of Eq. (2) involves two key approximations: (i) estimating the ACF from sampled data and (ii) truncating the integration bounds to a finite domain $[-\Delta_{\text{cut}},\Delta_{\text{cut}}]$. While collecting more data can systematically reduce the variance of the ACF estimate, addressing the second issue is more challenging. 19 Although extending the integration domain reduces bias due to truncation, it also increases variance for two reasons. First, integrating over a larger domain introduces more uncertain terms in the quadrature. Second, the variance of the integrand, the sampling ACF, grows at larger time lags. 20, 21, 22 To achieve a favorable trade-off between bias and variance, numerous algorithms have been proposed, some of which reformulate the integral to analyze alternative quantities, such as the mean squared displacement (MSD).23 Existing algorithms can be broadly classified into the four categories shown in Table 1, roughly in chronological order: (i) methods that directly analyze the ACF, (ii) methods that analyze the MSD, (iii) methods that apply cepstral analysis and (iv) spectral methods that directly analyze the power spectral density (PSD) of $\hat{x}(t)$. Section S1 of the Supporting Information lists known software implementations in each category. We will summarize each category below with equations for continuous and infinite time. This simplifies the notation as these quantities are not affected by sampling limitations. However, we will also discuss the limitations when the algorithms are applied to a finite amount of data.
@@
Green–Kubo (GK) ACF $=c(\Delta_{t})=\displaystyle\Bigl{\langle}\hat{x}(t_{0})\,\hat{x}(t_{0}+\Delta_{t})\Bigr{\rangle}$
Einstein–Helfand (EH) MSD $=\displaystyle\Bigl{\langle}\bigl{|}\hat{y}(t_{0}+\Delta_{t})-\hat{y}(t_{0})\bigr{|}^{2}\Bigr{\rangle}$
Cepstral Cepstrum $=\mathcal{C}(\Delta_{t})=\displaystyle\int_{-\infty}^{\infty}\log C(f)\,e^{2\pi if\Delta_{t}}\,\mathrm{d}f$
Spectral PSD $=C(f)=\displaystyle\int_{-\infty}^{\infty}c(\Delta_{t})\,e^{-2\pi if\Delta_{t}}\,\mathrm{d}\Delta_{t}$
@@ 2.1 Statistical properties of sampling power spectra
$\hat{x}_{d}(t)=h\sum_{n=-\infty}^{\infty}\hat{x}(hn)\,\delta(t-hn)$ (10)
where $\delta(t)$ is the Dirac delta function and the step size $h$ ensures that the units are consistent. Then $\hat{x}_{d}(t)$ can be substituted into the Wiener–Khinchin theorem in Eq. (5) to obtain:
@@
[…]
@@
$\hat{I}_{k}$ $=\frac{Fh}{2NM}\sum_{m=1}^{M}|\hat{X}^{(m)}_{k}|^{2}\sim\operatorname{Gamma}\left(\frac{\nu_{k}}{2},\frac{2I_{k}}{\nu_{k}}\right)$ (15)
@@ 2.2 Power spectrum model
$I^{\text{model}(S)}(f,\mathbf{b})=\exp\left(\sum_{s\in S}b_{s}f^{s}\right)$ (17)
where $\mathbf{b}$ is the vector of model parameters, and $S$ is the set of polynomial degrees, which must include 00. In most cases, using $S=\{0,1,2\}$ is sufficient. However, for spectra with large statistical uncertainties (when $M$ is small), it is advisable to use fewer terms. For example, $S=\{0\}$ is suitable for modeling a white-noise spectrum, while $S=\{0,1\}$ can extract a meaningful trend from highly noisy spectra. If the spectrum is expected to have a zero derivative at the origin, then $S=\{0,2\}$ can be used. A demonstration of the model for different sets of polynomial degrees $S$ is shown in Section S2 of the Supplementary Information. More elaborate models can be employed to extract additional properties from the spectrum, such as the exponential correlation time, which will be explored in a future paper. 65 Regardless of the model used, STACIE is designed to provide reliable error estimates. Therefore, it is recommended to select the model that minimizes the predicted uncertainty.
@@ 2.3 Parameter estimation
The switching function is $1/2$ at the cutoff frequency $f_{\text{cut}}$ and the exponent $\beta$ controls its steepness. As $\beta\rightarrow\infty$, the function approaches a hard cutoff. Using a smooth cutoff makes the search for suitable cutoff frequencies less sensitive to noise in the rescaled sampling PSD $\hat{I}_{k}$, as described in the next subsection. By default, STACIE uses $\beta=8$, for which the weight decreases from 0.9 to 0.1 over a frequency range of approximately $[0.76f_{\text{cut}},1.31f_{\text{cut}}]$. STACIE’s results are not very sensitive to the value of $\beta$. Although other switching functions exist, this form is simple and has conceptual advantages that become clearer when rewritten in terms of the hyperbolic tangent function:
@@
[…]
@@ 2.4 Frequency cutoff
$\operatorname{criterion}^{\text{CV2L}}_{j}=-\ln\mathcal{L}^{\text{CV2L}}\left(\hat{\mathbf{d}}^{(j)},\hat{\mathbf{C}}^{(j)}_{\mathbf{d}}\right)=\frac{P}{2}\ln(2\pi)+\frac{1}{2}\ln\left|\hat{\mathbf{C}}^{(j)}_{\mathbf{d}}\right|+\frac{1}{2}\bigl{(}\hat{\mathbf{d}}^{(j)}\bigr{)}^{T}\bigl{(}\hat{\mathbf{C}}^{(j)}_{\mathbf{d}}\bigr{)}^{-1}\hat{\mathbf{d}}^{(j)}$ (28)
@@
$\begin{split}\hat{\mathbf{b}}&=\sum_{j=1}^{J}W_{j}\,\hat{\mathbf{b}}^{(j)}\\ \hat{C}_{\mathbf{b},\mathbf{b}}&=\sum_{j=1}^{J}W_{j}\,\left(\hat{C}_{\mathbf{b}^{(j)},\mathbf{b}^{(j)}}+(\hat{\mathbf{b}}-\hat{\mathbf{b}}^{(j)})^{2}\right)\end{split}$ (30)
@@ 2.5 Summary of algorithmic hyperparameters
$\beta$ Steepness of the switching function $8$ $[8,20]$ (18)
$N_{\text{eff,min}}$ Start cutoff grid $5$ $[5,40]$ (24)
@@ 3.1 Minimal example
where $\hat{z}_{n}$ is drawn from a standard normal distribution. The parameter $\xi$ controls the magnitude of the noise term and $\phi$ controls the decay rate of the chain, with $10<\phi<$ 1. The discrete ACF of this chain decays exponentially:
@@
[…]
@@
Figure 3(b) displays the screen output of the Python script. First, STACIE prints the progress of cutoff frequency scan with three columns: $N_{\text{eff}}$, $-\ln\mathcal{L}^{\text{CV2L}}$ and $f_{\text{cut}}$. After completing the scan, all results are marginalized over the cutoff frequency and a summary of the results is printed on screen. Figures 3(c) and 3(d) are the plots saved to “minimal.pdf”.
@@ 3.2 Ionic electrical conductivity of an electrolyte
[…]
@@
[…]
@@
Figure 5: STACIE results for the ionic conductivity as a function of simulation time. (a) Ionic conductivity with its standard error estimated from five EMD trajectories. The value from Gullbrekken et al. 75 is shown as a black dot (slightly shifted to the right for clarity). STACIE results that pass all sanity checks (i.e. $N_{\text{eff}}>20P$ and Z-scores $>2$) are shown as filled dots. (b) Relative error of the estimated conductivity. (c) Effective number of grid points $N_{\text{eff}}$ used in the fit. The dashed line represents the recommended $N_{\text{eff}}=20\,P=40$. (d) The Z-scores derived from the regression cost function (green) and the cutoff criterion (yellow).
@@ 4 Validation
Using these three models, covariance kernels were either a single kernel model with specific parameters or a linear combination of two, as shown in Table 4. While the kernels were designed to produce diverse time series, they do share some common traits. In all cases, the theoretical PSD at zero frequency is exactly one, which is also the expected value of the autocorrelation integral (from $\infty$- ∞ to +$\infty$+ ∞) with $F=2$. Furthermore, for each test case, it was numerically verified that the first 20 amplitudes of the discrete PSD of the analytical kernel deviate by no more than 2.5% RMS from an even quadratic function, and the first 40 amplitudes deviate by no more than 10%. This ensures that the sampled sequences are all sufficiently long to resolve the spectrum of the slowest variations.
@@
Figure 7: Scaling of the errors of STACIE’s predictions for kernel exp1p in Table 4. Each data point represents a statistic over 64 independent tests for a given sequence length ($N$) and number of sequences averaged over ($M$, horizontal axis). Filled squares: the standard deviation on the predicted autocorrelation integral. Dotted lines: the predicted uncertainty. The slope of the light gray grid lines corresponds to the ideal decrease in error with increasing amount of data ($\propto 1/\sqrt{M}$) . Their spacing corresponds to the ideal decrease in error when quadrupling the sequence length.
@@
Figure 8: Assessment of the uncertainty quantification of STACIE for the kernel exp1p. Each data point represents a statistics over 64 independent tests for a given sequence length ($N$) and number of sequences averaged over ($M$, horizontal axis). Filled squares: the standard deviation of the predicted autocorrelation integral divided by the predicted uncertainty. Dots: the mean error divided by the predicted uncertainty.
@@ Acknowledgments
This research was funded by Ghent University with grant number BOF.STA.2017.0030.01. Computational resources were provided by Ghent University’s Stevin High-Performance Computing (HPC) infrastructure and the Flemish Supercomputer Center (VSC), funded by the Research Foundation Flanders (FWO). We extend our gratitude to Øystein Gullbrekken and Sondre Kvalvåg Schnell for their insightful responses to our questions regarding their electrolyte simulations.

### v2

@@ 1 Introduction
where $\hat{x}(t)$ is the time-correlated data, $t_{0}$ is the time origin, $\Delta_{t}$ is the time lag, and $\left\langle\cdot\right\rangle$ denotes an ensemble average. In this work, all stochastic quantities are denoted with a hat to distinguish them from deterministic ones, and the adjective “sampling” is used to distinguish them from expectation values in the text when needed for clarity. Practically, the average is taken over multiple time origins and independent sequences, e.g. different Cartesian components and/or different simulation runs. Examples of time-correlated inputs include particle velocities (used for diffusivity) and stress components (used for shear viscosity). This expression for the ACF assumes that $\langle\hat{x}\rangle=0$, which is often the case for transport properties. An exception is the computation of the bulk viscosity, where $\hat{x}(t)$ corresponds to the isotropic instantaneous pressure as a function of time, which fluctuates around its mean value. In this case, the mean pressure must be subtracted from $\hat{x}(t)$ before computing the ACF. 17
@@
where $F$ is a property-dependent factor. For example, for diffusivity, $F=1$, and for viscosity, $F=V/k_{\mathrm{B}}T$, where $V$ is the volume of the simulation cell, $T$ is the temperature, and $k_{\mathrm{B}}$ is the Boltzmann constant. 18, 3 Note that while it is more common to integrate from 0 to $\infty$, the integral is symmetric, allowing the limits to be extended to -$\infty$ to $\infty$ by including the factor $\frac{1}{2}$. This notation simplifies the discussion of STACIE in later sections. The integration of the ACF also yields the integrated correlation time:
@@
There are several sources of error when numerically integrating the ACF. In the context of MD simulations these include approximations of the function $\hat{x}(t)$, due to force fields, finite-size effects, Ewald convergence, etc. Even if these are well controlled, $\hat{x}(t)$ is a stochastic quantity (due to thermal fluctuations) and can only be sampled at a finite number of time steps (due to limited computational resources). As a result, a direct computation of $\mathcal{I}$ via numerical quadrature of Eq. (2) involves two key approximations: (i) estimating the ACF from sampled data and (ii) truncating the integration bounds to a finite domain $\Delta_{\text{cut}},\Delta_{\text{cut}}]$. While collecting more data can systematically reduce the variance of the ACF estimate, addressing the second issue is more challenging. 19 Although extending the integration domain reduces bias due to truncation, it also increases variance for two reasons. First, integrating over a larger domain introduces more uncertain terms in the quadrature. Second, the variance of the integrand, the sampling ACF, grows at larger time lags. 20, 21, 22 To achieve a favorable trade-off between bias and variance, numerous algorithms have been proposed, some of which reformulate the integral to analyze alternative quantities, such as the mean squared displacement (MSD).23 Existing algorithms can be broadly classified into the four categories shown in Table 1, roughly in chronological order: (i) methods that directly analyze the ACF, (ii) methods that analyze the MSD, (iii) methods that apply cepstral analysis and (iv) spectral methods that directly analyze the power spectral density (PSD) of $\hat{x}(t)$. Section S1 of the Supporting Information lists known software implementations in each category. We will summarize each category below with equations for continuous and infinite time. This simplifies the notation as these quantities are not affected by sampling limitations. However, we will also discuss the limitations when the algorithms are applied to a finite amount of data.
@@
Green–Kubo (GK) ACF $c(\Delta_{t})=\displaystyle\Bigl\langle\hat{x}(t_{0})\,\hat{x}(t_{0}+\Delta_{t})\Bigr\rangle$
Einstein–Helfand (EH) MSD $\Bigl\langle\bigl|\hat{y}(t_{0}+\Delta_{t})-\hat{y}(t_{0})\bigr|^{2}\Bigr\rangle$
Cepstral Cepstrum $\mathcal{C}(\Delta_{t})=\displaystyle\int_{-\infty}^{\infty}\log C(f)\,e^{2\pi if\Delta_{t}}\,\mathrm{d}f$
Spectral PSD $C(f)=\displaystyle\int_{-\infty}^{\infty}c(\Delta_{t})\,e^{-2\pi if\Delta_{t}}\,\mathrm{d}\Delta_{t}$
@@ 2.1 Statistical properties of sampling power spectra
$\hat{x}_{\text{d}}(t)=h\sum_{n=-\infty}^{\infty}\hat{x}(hn)\,\delta(t-hn)$ (10)
where $\delta(t)$ is the Dirac delta function and the step size $h$ ensures that the units are consistent. Then $\hat{x}_{\text{d}}(t)$ can be substituted into the Wiener–Khinchin theorem in Eq. (5) to obtain:
@@
The spectrum of the discrete-time process, $C_{\text{d}}(f)$, is periodic with period 1/h1/h64 Hence, if the domain of the PSD of the original continuous process, $C(f)$, is contained within [1/2h,1/2h][-1/2h,1/2h], it will be well reproduced by the spectrum of the discrete process, $C_{\text{d}}(f)$, without so-called aliasing artifacts. 64 Since the time step in MD simulations must be small enough to ensure that the Nyquist frequency, $1/2h$, is significantly larger than the highest frequency of the molecular system, there is no risk for aliasing when using all the time steps as input. In this case, we can safely use $\mathcal{I}\approx FC_{\text{d}}(0)/2$. However, writing results to disk at every time step for post-processing is not always possible due to storage limitations, for which a practical solution will be provided in Section 2.7.
@@
$\hat{I}_{k}$ $=\frac{Fh}{2NM}\sum_{m=1}^{M}\left|\hat{X}^{(m)}_{k}\right|^{2}\sim\operatorname{Gamma}\left(\frac{\nu_{k}}{2},\frac{2I_{k}}{\nu_{k}}\right)$ (15)
@@ 2.2 Power spectrum model
$I^{\text{model}}(f,\mathbf{b})=\exp\left(\sum_{s\in S}b_{s}f^{s}\right)$ (17)
where $\mathbf{b}$ is the vector of model parameters, and $S$ is the set of polynomial degrees, which must include 0. In most cases, using $S=\{0,1,2\}$ is sufficient. However, for spectra with large statistical uncertainties (when $M$ is small), it is advisable to use fewer terms. For example, $S=\{0\}$ is suitable for modeling a white-noise spectrum, while $S=\{0,1\}$ can extract a meaningful trend from highly noisy spectra. If the spectrum is expected to have a zero derivative at the origin, then $S=\{0,2\}$ can be used. A demonstration of the model for different sets of polynomial degrees $S$ is shown in Section S2 of the Supplementary Information. More elaborate models can be employed to extract additional properties from the spectrum, such as the exponential correlation time, which will be explored in a future paper. 65 Regardless of the model used, STACIE is designed to provide reliable error estimates. Therefore, it is recommended to select the model that minimizes the predicted uncertainty.
@@ 2.3 Parameter estimation
The switching function is $1/2$ at the cutoff frequency $f_{\text{cut}}$ and the exponent $\beta$ controls its steepness. As $\beta\rightarrow\infty$, the function approaches a hard cutoff. Using a smooth cutoff makes the search for suitable cutoff frequencies less sensitive to noise in the rescaled sampling PSD $\hat{I}_{k}$, as described in the next subsection. By default, STACIE uses $\beta=8$, for which the weight decreases from 0.9 to 0.1 over a frequency range of approximately $0.76f_{\text{cut}},1.31f_{\text{cut}}]$. STACIE’s results are not very sensitive to the value of $\beta$. Although other switching functions exist, this form is simple and has conceptual advantages that become clearer when rewritten in terms of the hyperbolic tangent function:
@@
$\begin{aligned} \hat{\mathcal{I}}&=\exp\left(\hat{b}_{0}+\frac{\hat{\sigma}^{2}_{b_{0}}}{2}\right)\\ \hat{\sigma}^{2}_{\mathcal{I}}&=\exp\Bigl(2\hat{b}_{0}+\hat{\sigma}^{2}_{b_{0}}\Bigr)\Bigl(\exp(\hat{\sigma}^{2}_{b_{0}})-1\Bigr)\end{aligned}$ (21)
@@ 2.4 Frequency cutoff
$\operatorname{criterion}^{\text{CV2L}}_{j}=-\ln\mathcal{L}^{\text{CV2L}}\left(\hat{\mathbf{d}}^{(j)},\hat{\mathbf{C}}^{(j)}_{\mathbf{d}}\right)=\frac{P}{2}\ln(2\pi)+\frac{1}{2}\ln\left|\hat{\mathbf{C}}^{(j)}_{\mathbf{d}}\right|+\frac{1}{2}\bigl(\hat{\mathbf{d}}^{(j)}\bigr)^{\top}\bigl(\hat{\mathbf{C}}^{(j)}_{\mathbf{d}}\bigr)^{-1}\hat{\mathbf{d}}^{(j)}$ (28)
@@
$\begin{split}\hat{\mathbf{b}}&=\sum_{j=1}^{J}W_{j}\,\hat{\mathbf{b}}^{(j)}\\ \hat{C}_{\mathbf{b},\mathbf{b}}&=\sum_{j=1}^{J}W_{j}\,\left(\hat{C}_{\mathbf{b}^{(j)},\mathbf{b}^{(j)}}+(\hat{\mathbf{b}}-\hat{\mathbf{b}}^{(j)})(\hat{\mathbf{b}}-\hat{\mathbf{b}}^{(j)})^{\top}\right)\end{split}$ (30)
@@ 2.5 Summary of algorithmic hyperparameters
$\beta$ Steepness of the switching function 88 $[8,20]$ (18)
$N_{\text{eff,min}}$ Start cutoff grid 55 $[5,40]$ (24)
@@ 3.1 Minimal example
where $\hat{z}_{n}$ is drawn from a standard normal distribution. The parameter $\xi$ controls the magnitude of the noise term and $\phi$ controls the decay rate of the chain, with $0<\phi<1$. The discrete ACF of this chain decays exponentially:
@@
For the minimal example, we used parameters $\phi=\frac{31}{33}$ and $\xi=\sqrt{\frac{8}{1089}}$, which are chosen to yield $\mathcal{I}=1$ and $\tau_{\text{int}}=16$. Because the ACF decays exponentially, we expect the PSD to have a Lorentzian shape with a maximum at the origin. Hence, we used the model in Eq. (17) with degrees $S=\{0,2\}$ to fit the low-frequency part of the spectrum, meaning that $P=2$. For this example, we aimed to estimate the autocorrelation integral with an error of about 2%, for which the suggested number of input time series, see Eq. (33), is $M\approx 1/(20\,P\,0.02^{2})\approx 64$.
@@
Figure 3(b) displays the screen output of the Python script. First, STACIE prints the progress of the cutoff frequency scan with three columns: $N_{\text{eff}}$, $\ln\mathcal{L}^{\text{CV2L}}$ and $f_{\text{cut}}$. After completing the scan, all results are marginalized over the cutoff frequency and a summary of the results is printed on screen. Figures 3(c) and 3(d) are the plots saved to “minimal.pdf”.
@@ 3.2 Ionic electrical conductivity of an electrolyte
where $\hat{J}^{\text{c}}_{i}(t)$ represents a Cartesian component $i$ of the charge current and the angle brackets $\langle\cdot\rangle$ denote the ensemble average. Practically, we average over the three Cartesian components, the time origins $t_{0}$, and the replicas of MD simulations with consistent settings. The factor $F$ is given by:
@@
where $V$ is the simulation box volume, $k_{\text{B}}$ is the Boltzmann constant, and $T$ is the temperature. Note that the usual factor $1/3$ is not needed in $F$ because in STACIE it is implied in the average over Cartesian components in $\langle\cdot\rangle$.
@@
Figure 5: STACIE results for the ionic conductivity as a function of simulation time. (a) Ionic conductivity with its standard error estimated from five EMD trajectories. The value from Gullbrekken et al. 75 is shown as a black dot (slightly shifted to the right for clarity). STACIE results that pass all sanity checks (i.e. $N_{\text{eff}}>20P$ and Z-scores >2>2) are shown as filled dots. (b) Relative error of the estimated conductivity. (c) Effective number of grid points $N_{\text{eff}}$ used in the fit. The dashed line represents the recommended $N_{\text{eff}}=20\,P=40$. (d) The Z-scores derived from the regression cost function (green) and the cutoff criterion (yellow).
@@ 4 Validation
Using these three models, covariance kernels were either a single kernel model with specific parameters or a linear combination of two, as shown in Table 4. While the kernels were designed to produce diverse time series, they do share some common traits. In all cases, the theoretical PSD at zero frequency is exactly one, which is also the expected value of the autocorrelation integral (from -$\infty$ to $\infty$) with $F=2$. Furthermore, for each test case, it was numerically verified that the first 20 amplitudes of the discrete PSD of the analytical kernel deviate by no more than 2.5% RMS from an even quadratic function, and the first 40 amplitudes deviate by no more than 10%. This ensures that the sampled sequences are all sufficiently long to resolve the spectrum of the slowest variations.
@@
Figure 7: Scaling of the errors of STACIE’s predictions for kernel exp1p in Table 4. Each data point represents a statistic over 64 independent tests for a given sequence length (NN) and number of sequences averaged over ($M$, horizontal axis). Filled squares: the standard deviation on the predicted autocorrelation integral. Dotted lines: the predicted uncertainty. The slope of the light gray grid lines corresponds to the ideal decrease in error with increasing amount of data ($\propto 1/\sqrt{M}$) . Their spacing corresponds to the ideal decrease in error when quadrupling the sequence length.
@@
Figure 8: Assessment of the uncertainty quantification of STACIE for the kernel exp1p. Each data point represents a statistics over 64 independent tests for a given sequence length (NN) and number of sequences averaged over ($M$, horizontal axis). Filled squares: the standard deviation of the predicted autocorrelation integral divided by the predicted uncertainty. Dots: the mean error divided by the predicted uncertainty.
@@ 5 Conclusions
Author Information
Author Contributions
G.T. performed and analyzed simulations, designed and validated algorithms and contributed to software development. D.F. and T.V. conceptualized and supervised the project. T.V. performed and analyzed simulations, designed and validated algorithms, and led software development. All authors wrote the original draft of the manuscript.
Notes
The authors declare no competing financial interest.
@@ Acknowledgments
This research was funded by the Research Board of Ghent University with grant number BOF/24J/2021/118. Computational resources were provided by Ghent University’s Stevin High-Performance Computing (HPC) infrastructure and the Flemish Supercomputer Center (VSC), funded by the Research Foundation Flanders (FWO). We extend our gratitude to Øystein Gullbrekken and Sondre Kvalvåg Schnell for their insightful responses to our questions regarding their electrolyte simulations.